OpenStudy (anonymous):

Help Please Linear Fx/Perpendicular (Medal/Fan) Find an equation for the linear function g(x) which is perpendicular to the line 7x-6y=12 and intersects the line 7x-6y=12 at x=36.

OpenStudy (anonymous):

So, y=(7/6)x-2

OpenStudy (anonymous):

The perpendicular would be y=(-6/7) ....I need help finding the y-intercept

OpenStudy (anonymous):

at \(x=36\) what is \(y\)?

OpenStudy (anonymous):

40?

OpenStudy (anonymous):

ok so find the equation of the line through the point \((16,40)\) with slope \(-\frac{6}{7}\)

OpenStudy (anonymous):

ok, ill work on in real quick

OpenStudy (anonymous):

jeez why did i say \((16,40)\) doh it should be \((36,40)\)

OpenStudy (anonymous):

i dont think this is right but i got y=-6/7x+(376/7)

OpenStudy (anonymous):

OK LOL

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

don't do what you did!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[y-40=-\frac{6}{7}(x-36)\] don't multiply out , pout in "standard form"

OpenStudy (anonymous):

\[7y-280=-6(x-36)\] etc

OpenStudy (anonymous):

never thought of it like that, ill try it, thanks. 1 sec

OpenStudy (anonymous):

got y=-6/7x+496/7

OpenStudy (anonymous):

thats right, im sure. thank you so much!

OpenStudy (anonymous):

if you insist

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i thought you wanted standard form

OpenStudy (anonymous):

nope, thanks thought you helped a lot

OpenStudy (anonymous):

oh nvm you wanted a "function"

OpenStudy (anonymous):

yw